The molecules A and B are made of C, H and O. The molecules were combusted and the combustion product was analyzed. Use the information below to determine the empirical formula for each molecule.

Acid

Mass Combusted (g)

Mass C (g)

Mass H (g)

Mass of O (g)

A

2.233

0.914

0.1015

1.218

B

1.564

0.9878

0.0822

0.4944
Answer #1

A)
we have mass of each elements as:
C: 0.914 g
H: 0.1015 g
O: 1.218 g


Divide by molar mass to get number of moles of each:
C: 0.914/12.01 = 0.0761
H: 0.1015/1.008 = 0.1007
O: 1.218/16.0 = 0.0761


Divide by smallest
C: 0.0761/0.0761 = 1
H: 0.1007/0.0761 = 4/3
O: 0.0761/0.0761 = 1

Multiply by 3 to get simplest whole number ratio:
C: 1*3 = 3
H: 4/3*3 = 4
O: 1*3 = 3


So empirical formula is:C3H4O3
Answer: C3H4O3

B)
we have mass of each elements as:
C: 0.9878 g
H: 0.0822 g
O: 0.4944 g


Divide by molar mass to get number of moles of each:
C: 0.9878/12.01 = 0.0822
H: 0.0822/1.008 = 0.0815
O: 0.4944/16.0 = 0.0309


Divide by smallest:
C: 0.0822/0.0309 = 8/3
H: 0.0815/0.0309 = 8/3
O: 0.0309/0.0309 = 1

Multiply by 3 to get simplest whole number ratio:
C: 8/3*3 = 8
H: 8/3*3 = 8
O: 1*3 = 3
So empirical formula is:C8H8O3
Answer: C8H8O3